OpenStudy (anonymous):

Hi All, I have been struggling with my first attempt at algebra and have done fine until now. We are doing Applications/Problem Solving of linear equations. My question is how to go about doing an application on a rectangle that gives the width as 44.6m and the Perimeter is 127m?

OpenStudy (saifoo.khan):

Use the perimeter formula, Perimeter = 2(L+B) insert the values and you are done!

OpenStudy (anonymous):

44.6*2=89.6 then 127-89.6=37.4 and divide by 2 to give 18.7 per length side

OpenStudy (anonymous):

Ok but I think it is 18.9?

OpenStudy (anonymous):

you are right, sorry i was doing it all in my head :)

OpenStudy (anonymous):

Thanks you guys you are really helpful though. I think I can figure how to compose the report itself if I can figure the darn equation structure.

OpenStudy (anonymous):

Wow, can I borrow your head, LOL!!!

OpenStudy (saifoo.khan):

Haha, dont forget to charge rent @ncjohnson6943

OpenStudy (anonymous):

haha it takes a lot of years (and a bachelors in mathematics) to do it all in your head. i may not wish all the math i have done upon my worst enemy :)

OpenStudy (anonymous):

LOL;) Thank you so much, I will use your equation to figure the rest of the info the instructor wants. Nice to find this place and who knows, maybe I can help someone one day:) Denise

OpenStudy (anonymous):

Thank you both, Saifoo and nc:)

OpenStudy (saifoo.khan):

No problemo.

OpenStudy (anonymous):

p=2l+2w so the equation would be 127=2L+2(44.6) so the answer would be L=3.9

OpenStudy (anonymous):

I get confused because how can the length be shorter than the width of the rectangle? If I have a width of 44.6m and the perimeter is 127m(I drew a rectangle to try and help me form my equation)than I see how we found the length with NC's figures but do you see what I am saying about the width being wider than the length? Maybe I have just lost my one, remaining marble:(

OpenStudy (anonymous):

it just means it is wider than it is longer, length doesnt have to be the longest dimension.

OpenStudy (anonymous):

Hmm, well that is important for me to learn;) Thank you MUCH!!!

OpenStudy (anonymous):

that is a good question it is wierd

OpenStudy (anonymous):

Ok, so P=2(L)+2(W), got it.

OpenStudy (radar):

Yes, that is the correct formula.

OpenStudy (anonymous):

Thanks you smartie pants;) I truly appreciate your time and patience with me, Denise

OpenStudy (anonymous):

Is that 3.9 the right answer? I can't figure how she got that?

OpenStudy (anonymous):

3.9 twice + 44.6 twice doesn't equal 127?

OpenStudy (radar):

127=2(44.6+L) 127/2=44.6+L 63.5=44.6 +L L=63.5-44.6=18.9

OpenStudy (radar):

18.9+18.9+44.6+44.6=127

OpenStudy (radar):

@nwcountry do you follow?

OpenStudy (anonymous):

That was what I got, I mean I didn't have the equation written right but I figured the answer using logic. Yes, 18.9 is the length correct?

OpenStudy (anonymous):

That 3.9 threw me though.

OpenStudy (radar):

yes, babygirl81 was the one that got 3.9

OpenStudy (anonymous):

Ok;) I will have to do a profile here and come back. What a neat site. I usually work alone because I can't find anyone at school that wants to study or practice Math ;)

OpenStudy (radar):

She had the correct formula, but must slipped up on the calculation

OpenStudy (anonymous):

I'm 59 and back in college, LOL!! But I love it and Math is my fave. I hated school when I was a kid, go figure;)

OpenStudy (radar):

Welcome and enjoy.

OpenStudy (anonymous):

Will do:) ttyl, Denise

OpenStudy (radar):

73 here lol

OpenStudy (anonymous):

Great! Nice to meet you Radar;) See you later, Denise

OpenStudy (anonymous):

Ok, I figured out how to present the equation the way my instructor wants it. Thought I would share it with you guys. Feedback welcome. I couldn't have gotten this far without your help;) 2(44.6)+2(L)=127 Perimeter, 89.2 + 2(L)=127 P, then I need to subtract 89.2 from each side leaving 2L =37.8, then I have to divide each side by 2 and I get my L=18.9 I suppose there are lots of ways of doing equations but I started recognizing how he's taught us. Thanks again, now all I have to do is put in the wordy part of the Problem Solving/Application which I can handle that:) bye for now, denise

OpenStudy (saifoo.khan):

Is that math 117 or 116?

OpenStudy (anonymous):

Math 60 it's called.

OpenStudy (saifoo.khan):

Introductory algebra! :D

OpenStudy (anonymous):

Yes:) The book is called Introductory & Intermediate Algebra Blitzer, heavier then my backpack, LOL!!! It will take me through 2 or 3 more terms though, same book.

OpenStudy (saifoo.khan):

lol, it has 1000+ pages! D:

OpenStudy (anonymous):

Yes, it does for sure;)

OpenStudy (anonymous):

Fun fun fun, LOL

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

This is nuts, I have (x+4)(x+6) and I keep getting x^2+10x+24. But the answer is x^2+11.+30 in the book? This is a practice question since they have the answers to odd #'d problems but not the even ones. We have to turn in the even ones.

OpenStudy (anonymous):

I have all my notes but he didn't do one like this in class:( I think he hates us, LOL!!!

OpenStudy (anonymous):

I'll be back, I need nutrition, LOL!!!

OpenStudy (saifoo.khan):

Your answer is correct ma'am.

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

you're joking?

OpenStudy (saifoo.khan):

No, really.

OpenStudy (anonymous):

I guess I better check the dang book answer again then.

OpenStudy (anonymous):

OMG, LOL, I must have looked at the wrong answer in the back, LOL!! Somebody beam me up and take me back to whatever planet I came from, LOL!!

OpenStudy (anonymous):

Thanks Saifoo, what a sweetie you are:) ttyl, Denise

OpenStudy (saifoo.khan):

LOLOL. Welcome ma'am.

OpenStudy (saifoo.khan):

And thanks